OpenStudy (anonymous):

Help on functions!! Intersection points between D(p)=500/p "AND" S(p)=p-10

OpenStudy (campbell_st):

you are looking for the point(s) where they functions are equal, so equate the functions and solve \[500/p = p - 10\] multiply through by p \[500 = p^2 - 10p\] now let it equal zero \[p^2 - 10p - 500 = 0\] and solve..for p... you might need to use the general quadratic equation, and there will be 2 points